How many grams of solid potassium fluoride should be added to 1.50 L of a 0.232 M hydrofluoric acid solution to prepare a buffer with a pH of 3.894 ?

grams potassium fluoride:

How many grams of solid potassium cyanide should be added to 0.500 L of a 0.131 M hydrocyanic acid solution to prepare a buffer with a pH of 8.738 ?

grams potassium cyanide:
Answer #1

Q1)

pK aof HF = 3.14

pH of buffer = 3.894

pH of buffer = pKa + log [KF]/[HF]

3.894 = 3.14 + log [KF]/[0.232]

thus [KF] = 1.3166M

mass of KF solid to be added to 1.5 L = 1.5L x molartiy x molar mass

= 1.5L x 1.3166mol/L x 58g/mol

=114.54 g

Q2)

pKa of HCN= 9.21

similar calculations

8.738 = 9.21 + log [KCN]/[0.131]

Thus [KCN] = 0.0442 M

and

mass of KCN to be added to 0.500L of solution = 0.5L x 0.0442mol/L x 65g/mol

=1.436 g